OpenStudy (anonymous):

How do I multiply functions together? Lets say we get the graph of f(x) and f(g). What would the graph of "fg" look like?

OpenStudy (amistre64):

same as you multiply normal numbers

OpenStudy (anonymous):

can you give examples?

OpenStudy (anonymous):

really depends on the functions doesn't it?

OpenStudy (amistre64):

x^3 +3x -2 5x -2 ---------------- 5x^4 +15x^2 -10x -2x^3 -6x +4 ------------------- 5x^4 -2x^3 +15x^2 -6x -10x +4

OpenStudy (anonymous):

f() = -1x g() = sin(x)

OpenStudy (anonymous):

i mean the graph part. you multiply by multiplying

OpenStudy (amistre64):

-x sin(x) then

OpenStudy (anonymous):

flips it over about the x axis

OpenStudy (anonymous):

oooh -x not -1 sorry

OpenStudy (anonymous):

sin(x) is stuck between 1 and -1 so -xsin(x) is stuck between the two lines y = x and y = -x

OpenStudy (amistre64):

the graph of the sin is flipped annd is bounded by y=x line right?

OpenStudy (anonymous):

very nice picture actually

OpenStudy (anonymous):

y = x and y = -x.

OpenStudy (anonymous):

Alright then. Lets say you were given only a graph of 2 functions. How would you find out how the graph of fg look like then?

OpenStudy (anonymous):

not so easy

OpenStudy (anonymous):

That's one of the questions on my practice SAT 2 Math 2. Ugh...

OpenStudy (anonymous):

just think of a simple example. you know what y = x looks like, you know what y = x^2 looks like, but from the two graphs how could you visualize y = x^3?

OpenStudy (amistre64):

if f(x) = 3; and g(x) = 10; then id say h(x) = 30 if f(x) = 4; and g(x) = 5; then id say h(x) = 20 if f(x) = 1/2; and g(x) = 38; then id say h(x) = 19

OpenStudy (amistre64):

you multiply the corresponding y values to get the resulting new value

OpenStudy (anonymous):

^ That made everything so much more simple. Thanks!

OpenStudy (amistre64):

:)